I have another question that I need assistance with. Any assistance is greatly appreciated. Thank you.
A plane took 1 hour longer to travel 560 miles on the first portion of a flight than it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for the second part of the trip?

Answer by stanbon(75887)

A plane took 1 hour longer to travel 560 miles on the first portion of a flight than it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for the second part of the trip?
-----------------
First portion DATA:
Distance = 560 miles; time = x+1 hrs ; Rate = d/t = 560/(x+1) hrs
----------------
Second portion DATA:
Distance = 480 miles ; time = x hrs ; Rate = d/t = 480/x hrs
--------------------
EQUATION:
rate = rate
560/(x+1) = 480/x
7/(x+1) = 6/x
7x = 6x+6
x = 6 hours (Time of second portion of flight)
===============
Cheers,
Stan H.